# A Novel Compact LLM Framework for Local, High-Privacy EHR Data Applications

### v1

@@ 4.3 Results
Time range Method Preprocessing status Classified correctly ($P_{c}$) Classified incorrectly ($P_{i}$) No information is provided ($P_{l}$)
@@
Table 1 represents the experiment results using our Private HNC Dataset. The results from our investigations substantiate the efficacy of the preprocessing steps integrated into our framework. The data clearly indicates that this preprocessing significantly enhances the classification accuracy across various modeling paradigms, encompassing both zero-shot and few-shot learning approaches. This enhancement is particularly crucial when employing compact, locally-operated LLMs for processing EHR datasets. Such models, due to their smaller size relative to their more extensive counterparts, often face computational and memory constraints that can limit their effectiveness.
By mitigating these challenges through our targeted preprocessing steps, we are able to considerably improve the utility of smaller LLMs in high-privacy, local computational environments, thereby enhancing their applicability and effectiveness in real-world clinical settings. This finding not only advances our understanding of model scalability and efficiency but also opens avenues for further research into optimizing LLM frameworks for specialized data-intensive tasks in constrained environments.
Method Preprocessing status
Results (Patient)
Results (Hospital admission)
@@
Table 2: LLM classification results for “so-called” sensitivity analysis for MIMIC-IV dataset with different preprocessing and shot settings. Results (Patient) denotes Percentage of patients with relevant ICD code & Been classified with metastasis by LLM out of all patients with relevant ICD code (#$\frac{\#(M_{LLM}^{i}\cap M_{ICD}^{i})}{\#(M_{ICD}^{i})}$). Results (Hospital admission) denotes percentage of hospital admissions with relevant ICD code & Been classified with metastasis by LLM out of all patients with relevant ICD code (#$\frac{\#(M_{LLM}^{ha}\cap M_{ICD}^{ha})}{\#(M_{ICD}^{ha})}$). The rows of 2b and 7b are for Gemma-2b and for Gemma-7b.
The results in Table 2 highlight the substantial impact of preprocessing on the performance of LLMs when applied to clinical notes from the MIMIC-IV dataset across both zero-shot and few-shot learning settings.
Preprocessing leads to significant improvements in classification accuracy across various learning settings. In zero-shot learning on a per-patient basis, preprocessing results in a substantial increase in accuracy compared to models without preprocessing. Similarly, both few-shot settings show large performance gains when preprocessing is applied. This pattern holds true in evaluations based on hospital admissions, where the preprocessed zero-shot model performs far better than the model without preprocessing. The same trend is observed in few-shot models, with preprocessing consistently yielding significantly higher accuracy than models that do not use preprocessing.
Another notable trend, similar to our findings with the private dataset, is that the model’s performance tends to decline as the number of provided examples increases. And it is likely because longer prompts with more examples make it difficult for relatively small models to focus on the key information in the clinical notes.
We also observe that with a limited number of samples, the combination of preprocessing and zero-shot learning outperforms fine-tuning the model on those samples before classification. This suggests that preprocessing can be more effective in guiding the model to relevant information than relying on a small amount of fine-tuning data.
Table 2 focuses solely on the “so-called” sensitivity (i.e., power) of the method. Additional Table 3 in the Appendix provide further validation by assessing the “so-called” specificity. This table shows that all methods perform well in terms of specificity. As a result, the primary emphasis should be on sensitivity, where the preprocessed dataset outperforms the others.
Summary
All methods demonstrate high specificity and accurately identify non-metastatic cases, the key advantage lies in the models’ sensitivity. The overall conclusion from these results is that preprocessing, which extracts and highlights critical information from clinical notes, significantly improves LLM performance. This makes preprocessing an essential step when using smaller LLMs on unstructured and complex datasets like EHRs. However, as the number of examples (shots) increases, the performance tends to decline, probably due to input length limitations in smaller models. This highlights the importance of preprocessing for maintaining model efficiency in resource-constrained environments. Furthermore, the findings suggest that when ground truth data is scarce, combining preprocessing with zero-shot learning may be more effective than fine-tuning the model with a limited number of samples. This indicates that in situations where annotation is labor-intensive, particularly for EHR datasets, using preprocessing with zero-shot learning is a more efficient and accurate approach than manual annotation.
5 Discussion
In this study, we introduce a novel compact LLM framework tailored for local deployment on EHR data, addressing significant challenges such as privacy concerns and computational limitations in healthcare settings. Our approach, which integrates a preprocessing step such as regular expressions, significantly enhances the performance of compact LLMs. This preprocessing not only reduces the data’s complexity before it reaches the LLM but also ensures that the models focus on the most relevant information, thereby improving accuracy and efficiency.
Using metastasis as an example, we demonstrate the importance of preprocessing using the Private HNC dataset and MIMIC-IV dataset. The results of our experiments underscore the potential of compact LLMs to perform robustly even under constraints, aligning with previous research that emphasizes the scalability and adaptability of LLMs in various domains. However, unlike previous studies that primarily focus on maximizing model performance without regard for computational efficiency, our work specifically targets the balance between performance and resource utilization, which is critical in many practical healthcare applications.
Our study offers practical insights for the local deployment of LLMs, particularly when utilizing models of a smaller scale. We recommend prioritizing the preprocessing of datasets to enhance model performance rather than focusing primarily on fine-tuning. This approach facilitates more effective use of limited computational resources and maintains data privacy, proving especially advantageous in settings where extensive model training is impractical.
While our findings are promising, the study is not without limitations. The framework’s reliance on regex-based preprocessing may not capture all nuances of clinical language, potentially overlooking subtle but clinically relevant information not explicitly highlighted by the regex patterns. Additionally, the performance of our framework is still dependent on the quality and scope of the input data, which may vary across different healthcare settings.
Future research could explore the integration of more sophisticated NLP techniques, such as context-aware embeddings or deep learning-based entity recognition, to enhance the accuracy and adaptability of the preprocessing step. Further studies could also evaluate the framework’s performance across a broader range of diseases and clinical conditions to better understand its generalizability.
Appendix A Additional table for specificity analysis
Method Processing status
Results (Patient)
Results (Hospital admission)
@@ Appendix A Additional table for specificity analysis
Table 3: LLM classification results for “so-called” specificity analysis for MIMIC-IV dataset with different preprocessing and shot settings. Results (Patient) denotes Percentage of patients without relevant ICD code & Not been classified with metastasis by LLM out of all patients without relevant ICD code (#$\frac{\#(N_{LLM}^{i}\cap N_{ICD}^{i})}{\#(N_{ICD}^{i})}$). Results (Hospital admission) denotes percentage of hospital admissions without relevant ICD code & Not been classified with metastasis by LLM out of all patients without relevant ICD code (#$\frac{\#(N_{LLM}^{ha}\cap N_{ICD}^{ha})}{\#(N_{ICD}^{ha})}$).
@@ Appendix B Implementation details
Predictions using zero-shot or few-shot paradigms typically take 1 to 2 weeks due to the sequential processing of individual records. Fine-tuning the models on 100 samples is significantly faster, requiring less than 10 minutes, and subsequent classification of the entire dataset takes approximately 2 days.
The implementation code for the MIMIC-IV dataset is available publicly on https://anonymous.4open.science/r/LLM_EHR-175A.

### v2

@@ Appendix A Additional Tables
Time range Method Preprocessing Status Classified Correctly ($P_{c}$) Classified Incorrectly ($P_{i}$) No Information ($P_{l}$)
@@
Table 2: LLM classification results under different preprocessing and shot settings for the MIMIC-IV dataset.
Method Preprocessing Status Results (Patient) Results (Hospital Admission)
@@
Notes: Results (Patient) denotes percentage of patients with relevant ICD code & Been classified with metastasis by LLM out of all patients with relevant ICD code (#$\frac{\#(M_{LLM}^{i}\cap M_{ICD}^{i})}{\#(M_{ICD}^{i})}$). Results (Hospital admission) denotes percentage of hospital admissions with relevant ICD code & Been classified with metastasis by LLM out of all patients with relevant ICD code (#$\frac{\#(M_{LLM}^{ha}\cap M_{ICD}^{ha})}{\#(M_{ICD}^{ha})}$). The rows of 2B and 7B are for Gemma-2B and for Gemma-7B.
Table 3: LLM classification results for specificity analysis on the MIMIC-IV dataset.
Method Processing Status Results (Patient) Results (Hospital Admission)
@@
Notes: Results (Patient) denotes Percentage of patients without relevant ICD code & Not been classified with metastasis by LLM out of all patients without relevant ICD code (#$\frac{\#(N_{LLM}^{i}\cap N_{ICD}^{i})}{\#(N_{ICD}^{i})}$). Results (Hospital admission) denotes percentage of hospital admissions without relevant ICD code & Not been classified with metastasis by LLM out of all patients without relevant ICD code (#$\frac{\#(N_{LLM}^{ha}\cap N_{ICD}^{ha})}{\#(N_{ICD}^{ha})}$).
Table 4: Comparison of different LLMs under various shot settings on the MIMIC-IV dataset.
Shot Setting Processing State Gemma-7B-it LLaMA-2-7B-Chat-Med Bio-Medical-LLaMA-3-8B
Zero-shot Preprocessed with regex 0.835 (0.831-0.836) 0.749 (0.729-0.763) 0.595 (0.584-0.606)
No Preprocessing 0.080 (0.078-0.081) 0.384 (0.367-0.401) 0.495 (0.482-0.516)
Three-shot Preprocessed with regex 0.813 (0.809-0.814) 0.464 (0.438-0.494) 0.712 (0.699-0.719)
No Preprocessing 0.085 (0.073-0.091) 0.217 (0.203-0.235) 0.280 (0.271-0.291)
Six-shot Preprocessed with regex 0.616 (0.609-0.621) 0.618 (0.606-0.636) 0.546 (0.538-0.555)
No Preprocessing 0.0058 (0.0024-0.0073) 0.562 (0.536-0.580) 0.427 (0.413-0.452)
Notes: Results denote the percentage of patients with relevant ICD codes classified with metastasis by LLM out of patients with relevant ICD codes. Results are based on running each model 10 times on a 5% sample of the dataset. Values are reported as mean (range) across runs.
Table 5: Comparison of preprocessing approaches under different shot settings on MIMIC-IV dataset.
Shot Setting Preprocessing Status Mean of $P_{c}^{i}$ Range of $P_{c}^{i}$
Zero-shot Preprocessed with regex 0.8493 0.8356 - 0.8630
Preprocessed with RAG (Top 2) 0.6863 0.6712 - 0.6986
Preprocessed with RAG (Top 3) 0.6534 0.6301 - 0.6575
Preprocessed with RAG (Top 6) 0.6657 0.6438 - 0.6712
No preprocessing 0.0671 0.0548 - 0.0685
Three-shot Preprocessed with regex 0.7972 0.7945 - 0.8082
Preprocessed with RAG (Top 2) 0.7383 0.5342 - 0.7671
Preprocessed with RAG (Top 3) 0.6616 0.5342 - 0.7123
Preprocessed with RAG (Top 6) 0.4494 0.4384 - 0.4658
No preprocessing 0.10 0.0548 - 0.1233
Six-shot Preprocessed with regex 0.6685 0.6575 - 0.6849
Preprocessed with RAG (Top 2) 0.6438 0.6301 - 0.6575
Preprocessed with RAG (Top 3) 0.4055 0.3973 - 0.4110
Preprocessed with RAG (Top 6) 0.1630 0.1507 - 0.1644
No preprocessing 0.0247 0.0137 - 0.0274
Notes: Results denote the percentage of patients with relevant ICD codes classified with metastasis by LLM out of patients with relevant ICD codes. Results are based on running the Gemma-7B-it model 10 times on a 1% sample of the dataset. Values are reported as mean and range across runs.
@@ Appendix B Implementation Details
For Gemma-7B-it model, predictions under zero-shot or few-shot paradigms typically take 1 to 2 weeks due to the sequential processing of individual records across the entire dataset. Fine-tuning the models on 100 samples is significantly faster; it requires less than 10 minutes. The subsequent classification of the entire dataset takes approximately 2 days. The regex-based approach and the RAG-based approach (with top-2 sentence retrieval) have similar processing times while retrieving the top 6 sentences using RAG increases runtime by approximately 2.5 times. LLaMA models generally take longer than Gemma models, with Bio-Medical-LLaMA-3-8B taking roughly twice as long, but they require less GPU memory and can operate within a 16GB VRAM environment.